Write an equation of the line that is perpendicular to the given line and passes through the given point y=2x-1 (2,4)
Answer by Cromlix(4381)

Hi there,
y = 2x - 1
Gradient = 2
Lines that are perpendicular to
one another have gradients that
multiply together to give -1
m1 x m2 = -1
Using gradient = 2
2 x m2 = -1
m2 = -1/2
Using line equation:
y - b = m(x - a)
m = -1/2
(a,b) = (2,4)
y - 4 = -1/2(x - 2)
y - 4 = -1/2x + 1
y = -1/2x + 1 + 4
y = -1/2x + 5
or multiply through by 2
2y = -x + 10
Hope this helps :-)